Prove whether or not the program segment x≔3 z≔x-y+2 if y>0 then z≔z+3 else z≔2   ...

  1. Prove whether or not the program segment

x≔3

z≔x-y+2

if y>0 then

z≔z+3

else

z≔2

   is partially correct with respect to the initial assertion y=4 and the final assertion z=6
Answer #1


x = 3 Z = X-Y+2 if yyo then Z: - Z+3 else Z=2 For Y=4 x - 3 Z:= X - Y + 2 - 3- 4+2 - since Y=470 Z: - Z+3 . Z= 1+3 = 476 The